OpenStudy (anonymous):

*MEDAL* Solve. 2x +6y = -12

OpenStudy (anonymous):

2x +6y =-12 2x = -12 -6y x = (12-6y) ------- 2 substitute that statement for x 2(12-6y) +6y = -12 ---- 2 factor left side 24 - 12y +6y = -12 solve for y then solve for x

OpenStudy (anonymous):

24 -6y =-12 24 + 12 = 6y 36 = 6y 6 = y insert 6 for y and solve for x

OpenStudy (anonymous):

24 - 18y = -12 (add 24 to each side) -18y = -36 (divide -18 by each side) y = 2?

OpenStudy (anonymous):

What are you solving for?

OpenStudy (anonymous):

Both, i guess. both x and y.

OpenStudy (anonymous):

no add 18y to both sides ad 12 to both sides to get: 36 = 6y

OpenStudy (anonymous):

Hard to believe thats the question (solving for both)

OpenStudy (anonymous):

you cant add 24 to get rid of 24

OpenStudy (anonymous):

ooh, so now divide 6 by each side. So y = 6

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

insert 6 for y and solve for x

OpenStudy (anonymous):

So now it's 2x + 6(6) = -12?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

you got this

OpenStudy (anonymous):

2x + 36 = -12 -36 -36 ___________________ 2x = -48 2 2 x = -24 I feel like I did something wrong...

OpenStudy (anonymous):

looks good to me 2(-24) + 36 = -12 -48 + 36 = -12 -12 = -12

OpenStudy (campbell_st):

my question is what do you need to solve for.... which letter... my next question is... is there another equation....? you cannot solve 2 equation with 2 unknowns......

OpenStudy (anonymous):

Thanks :D

OpenStudy (anonymous):

You can use this technique to solve for three variables too, it's wold

OpenStudy (anonymous):

(wild)

OpenStudy (anonymous):

The original question is this: @campbell_st

OpenStudy (anonymous):

campbell, you can solve for 2 or 3 unknowns easily with the right tools

OpenStudy (campbell_st):

what happens if x = 0 and y = -2.... thats a solution it gives an answer of -12 along with x = -12 and y = 1 that gives an answer of -12

OpenStudy (campbell_st):

thats better, you are solving 2 equations in 2 unknowns so you should get a unqique solution

OpenStudy (anonymous):

that is A possible solution but not THE solution. Because when you solve for the first step: x = (12-6y) ------- 2 you would have gotten a different equation if it were your numbers

OpenStudy (anonymous):

Choices: (2, 1) (0, -2) (-2, 0) (1, 2)

OpenStudy (campbell_st):

multiply the 1st equation by 5 and the 2nd equation by 6, this will allow you to eliminate y. add the 2 new equations 10x + 30y = -60 30x - 30y = 60 ------------------ 40x = 0 x = 0 subsitute into either of the original equations to find y hope this helps

OpenStudy (anonymous):

it does. By ruling other options out (since other then 1 has x = 0), then it would be (0, -2). B.

OpenStudy (anonymous):

well well well, I wasn't shown the second equation that narrows it down and campbell is right Special you have to present all the info

OpenStudy (campbell_st):

@ehuman, I agree with the right skills or tools you can solve anything.... but you cannot solve 1 equation with 2 unknowns. All you can do is change the subject

OpenStudy (campbell_st):

thats correct well done

OpenStudy (campbell_st):

sorry, I just in year 10 maths.... and have to et back to my work

OpenStudy (anonymous):

Insert the first solved variable into the second equation to get the correct answer

OpenStudy (anonymous):

Sorry, @ehuman . My friend showed me this way of working this where if you figure out the first problem, it should rule out the other 3 options. Meh, didn't work with this one! Haha

OpenStudy (campbell_st):

actually substituting into either of the original equations will give the correct solution. You are really finding the point on the number plane where the 2 lines intersect.